OpenStudy (aimmebgood):

factor completely 10r^(2) - 1,250.

OpenStudy (anonymous):

Remember the difference of squares: \[a^2-b^2 =(a-b)(a+b) \]

OpenStudy (anonymous):

so \(a^2 = 10r^2\) and \(b^2=1,250\). So solve for \(a\) and \(b\). Then plug them back into \((a-b)(a+b)\).

OpenStudy (anonymous):

Does that not help?

OpenStudy (aimmebgood):

on confused on how to solve for a and b. Or how to start to solve for a and b.